# Whose Boat Does it Float? Improving Personalization in Preference Tuning via Inferred User Personas

### v1

@@ Abstract
LLMs are tuned to follow instructions (aligned) by learning which of two outputs users prefer for a prompt. However, this preference data format does not convey why users prefer responses that are chosen or rejected, so LLMs trained on these datasets cannot tailor responses to varied user needs. To surface these parameters of personalization, we apply abductive reasoning to preference data, inferring needs and interests of users, i.e. personas, that may prefer each output. We test this idea in two steps: Persona Inference (PI)—abductively inferring personas of users who prefer chosen or rejected outputs—and Persona Tailoring (PT)—training models to tailor responses to personas from PI. We find: 1) LLMs infer personas accurately explaining why different users may prefer both chosen or rejected outputs; 2) Training on preference data augmented with PI personas via PT boosts personalization, enabling models to support user-written personas; and 3) Rejected response personas form harder personalization evaluations, showing PT better aids users with uncommon preferences versus typical alignment methods. We argue for an abductive view of preferences for personalization, asking not only which response is better but when, why, and for whom. [^1] Code/data are publicly available at https://github.com/Pinafore/alignment-personalization
@@ 1 Every Preference Happens for a Reason
Current methods for aligning large language models (LLMs) predominantly rely on preference data Ji et al. (2023), which is created by finding: for an input prompt, which of two outputs do users prefer? LLMs are then trained to give responses like the majority chosen response and unlike the rejected one, improving LLM output quality in dialogue Köpf et al. (2024), question answering (Fan et al., 2019, QA), and summarization Völske et al. (2017).
While useful, preference datasets assume chosen responses are universally better, failing to consider why users prefer responses. In reality, some users may prefer rejected outputs, even if their reasons are less common. In Figure 2 (left), most users prefer the chosen output for taking brownies to a cake sale. While users valuing simplicity may prefer this direct response, practical users may prefer the rejected one which also has packaging logistics (Figure 2, middle). As users prefer responses for varied reasons Kirk et al. (2024b), models tuned on preference data should personalize outputs to meet these specific, individual needs Salemi et al. (2024).
A common, interpretable method for personalization involves users verbalizing their needs and interests as an extra input (e.g. “I like short answers”)—forming a system prompt persona to tailor model responses Zhang et al. (2024c). However, current preference data formats do not have personas conveying why users prefer responses, omitting signals for training personalized models Lee et al. (2024).
Figure 1: Alignment methods with typical preference datasets like DPO cannot fully cater to a user’s specified personas. To overcome this, we train models on preference data augmented with LLM-inferred personas, which we call persona tailoring.
@@
The only way for models tuned on existing preference data formats to use a persona is by adding it to the prompt in inference, hoping the model treats it as an extra instruction to follow Deb et al. (2022). However, typical preference tuning methods like direct preference optimization (Rafailov et al., 2024, DPO) lack this personalization. In Figure 1, DPO tells a sober user to hire a bartender (left) for their party, while a user desiring a small list gets 10 ideas (right); such responses are unhelpful, inappropriate, and fail to support the users’ specified personas.
As models’ poor personalization stems from a lack of training on personas, we propose abductive reasoning Peirce (1974) to augment preference training data with LLM-inferred personas. Abduction infers hidden contexts that explain a given outcome Zhao et al. (2024a). Similarly, we adopt this reasoning to infer (i.e. adduce) hidden needs, interests, and traits of users, i.e., personas Tseng et al. (2024), that explain why chosen and rejected outputs may be preferred. If LLMs surface valid personas of users who may prefer chosen or rejected responses, we can attach them to preference datasets and train models that tailor to user-specified needs, enhancing model personalization (Figure 1, blue).
@@
We first thoroughly test LLM abilities in PI on dialogue Bai et al. (2022a), QA Ji et al. (2024), and education Balepur et al. (2024b) preference data. LLM personas accurately explain why users may prefer chosen and rejected outputs; LLaMA-405B has 91% accuracy, judged by GPT-4o (section 3.1). Also, GPT-4o often judges chosen and rejected response personas as tied in quality (section 3.2), and qualitatively, rejected ones are rated as highly-plausible but applying to less users (section 3.3). Thus, users may prefer rejected outputs for uncommon but valid reasons. Personas also describe implicit differences Hoyle et al. (2023) in chosen and rejected responses; we find in BeaverTails Ji et al. (2024), users in chosen response personas are “meticulous”, while rejected ones are “direct”, showing the data annotators tend to prefer verbose responses Zheng et al. (2024).
As PI yields accurate personas, we augment preference data with PI personas using LLaMA-405B. We then train LLaMA-8B on this data for the reverse task of PT: using prompts and inferred personas as inputs to give tailored responses (Figure 2, right). We test three strategies: prompting (Brown et al., 2020, $\textsc{PT}_{\textsc{fs}}$), fine-tuning (Chung et al., 2024, $\textsc{PT}_{\textsc{sft}}$), and DPO (Rafailov et al., 2024, $\textsc{PT}_{\textsc{dpo}}$).
Training on LLM personas via PT largely aids personalization (section 5.1), with $\textsc{PT}_{\textsc{dpo}}$ as the strongest method (section 5.2). Further, DPO tuned on preference datasets without personas cannot always tailor to personas in inference; notably, $\textsc{PT}_{\textsc{dpo}}$ is judged as much stronger than DPO on uncommon but valid needs in the personas from rejected responses (58% average improvement in personalization), showing rejected responses form valuable, harder evaluations for personalization (section 5.3). Lastly, we have ten users write diverse personas and qualitatively assess $\textsc{PT}_{\textsc{dpo}}$ and DPO outputs (section 5.4). The same users find $\textsc{PT}_{\textsc{dpo}}$ better personalizes to their written needs, showing its utility to aid users downstream.
@@
2) We employ PI for data augmentation and release persona-augmented versions of question answering, dialogue, and education preference datasets.
3) We show persona tailoring—training on persona-augmented preference data—enhances personalization for LLM-inferred and user-specified needs.
@@ 2 Persona Inference Setup
The first step in our personalization method is inferring why users may prefer each output in standard preference data (Figure 2, middle). We adopt abduction Peirce (1974)—which explains given outcomes—to infer why preference data responses may be preferred via Persona Inference (PI):
@@
Following the taxonomy of Chen et al. (2024a), our personas are closest to demographic personas. However, we infer high-level traits (Figure 2, middle), like information needs, interests, and personality types, that alter how users want outputs shown, rather than protected attributes (e.g. race) to avoid stereotyping Kantharuban et al. (2024). This section gives our PI models (section 2.1) and datasets (section 2.2).
@@ 2.1 Models
We test nine LLMs in PI: Claude (Anthropic, 2023, Sonnet, Haiku, Opus), GPT (Achiam et al., 2023, 3.5, 4, 4o), and LLaMA-3.1 Instruct (Dubey et al., 2024, 8B, 70B, 405B). We use 0 temperature and five-shot prompts written by one author, using rules in Schulhoff et al. (2024). Exemplars use the format below, where highlights show generated text:
@@ 2.2 Datasets
1) BeaverTails Ji et al. (2024) has advice questions $p$ and candidate answers $\mathcal{R}$ on 14 harm categories including politics, privacy, and unethical behavior.
2) Stanford Human Preferences (Ethayarajh et al., 2022, SHP) has Reddit post questions $p$ on r/ask and advice forums with user-written answers $\mathcal{R}$.
@@
4) Mnemonic Balepur et al. (2024b) has vocabulary terms $p$ and keyword mnemonics $\mathcal{R}$. Mnemonics are study aids that help users learn $p$’s meaning.
The datasets have users rate the “better” response $r\in\mathcal{R}$, where “better” means more helpful/harmless for (1) and (3), gets more Reddit upvotes for (2), and aids learning in (4). The overall better $r$ is chosen ($r_{\texttt{C}}$) and the other is rejected ($r_{\texttt{R}}$). For each entry, we alter inputs $r_{1}$ and $r_{2}$ in PI to get chosen personas $\mathcal{P}_{\texttt{C}}=\textsc{PI}(p,r_{1}=r_{\texttt{C}},r_{2}=r_{\texttt{R}})$ and rejected personas [^2] We use “rejected” for brevity, not to imply they are worse. $\mathcal{P}_{\texttt{R}}=\textsc{PI}(p,r_{1}=r_{\texttt{R}},r_{2}=r_{\texttt{C}})$ for $r_{\texttt{C}}$ and $r_{\texttt{R}}$. As $r_{\texttt{R}}$ is preferred less often, we assume $\mathcal{P}_{\texttt{R}}$ has less common/popular needs (tested in section 3.3).
In BeaverTails and SHP, some responses $r_{\texttt{R}}$ are deliberately low-quality, with harmful or inaccurate content Liu et al. (2024b). These are out-of-scope, as all $\mathcal{P}_{\texttt{R}}$ will be undesirable [^3] Such as “the user is unethical” or “likes misleading users”. and we do not want to have models tailor to them (section 9). Thus, we use a split of outputs labeled “safe” in BeaverTails and outputs with 10+ upvotes in SHP. Post-filtering, we randomly sample 300 entries in each dataset to form 600 PI inputs (data details in Appendix A.1).
3 Understanding LLM-Inferred Personas
@@
Before directly using LLM personas to train personalized models (section 4), we first thoroughly assess their quality and differences in chosen and rejected personas, providing insights into the types of personas LLMs infer. This section finds LLMs’ chosen and rejected personas: 1) accurately explain why users may prefer each preference dataset response (section 3.1); 2) convey similarly-valid needs (section 3.2, section 3.3); and 3) reveal preference data trends (section 3.4), showing they are reliable signals for augmenting preference data.
@@ 3.1 LLMs Accurately Infer Personas
We first ensure LLM personas satisfy abduction’s goal: justifying why its input response is better (section 2). To test this, we have GPT-4o, a strong judge Kim et al. (2024), pick the best response out of $r_{\texttt{C}}$ and $r_{\texttt{R}}$ for a prompt and persona. A chosen persona $\mathcal{P_{\texttt{C}}}$ is accurate if GPT-4o picks $r_{\texttt{C}}$ over $r_{\texttt{R}}$ (same for $\mathcal{P_{\texttt{R}}}$). GPT-4o shows 86% agreement with one Ph.D. student (Appendix A.4), so the metric is reliable.
LLMs infer accurate $\mathcal{P_{\texttt{C}}}$ and $\mathcal{P_{\texttt{R}}}$; the judge picks the intended response in most cases (Figure 3). $\mathcal{P_{\texttt{R}}}$ is always less accurate, consistent with work that shows LLMs struggle to justify incorrect answers Michael et al. (2023); Balepur et al. (2024a). However, some models have small accuracy gaps (0.06 for L-405B), so while $\mathcal{P_{\texttt{R}}}$ is harder to infer, it can still identify plausible needs of users. Finally, Mnemonic accuracy is lowest, as LLMs must infer why outputs aid learning, which is likely harder than why they are helpful or harmless. Thus, in specific, subjective domains Padmakumar et al. (2024), it may be useful to directly elicit why annotators prefer responses in preference data collection, rather than relying on LLMs to infer them.
@@ 3.2 LLMs Judge Personas as Similar Quality
As LLMs infer accurate chosen/rejected personas $\mathcal{P}_{\texttt{C}}$/$\mathcal{P}_{\texttt{R}}$, we now compare $\mathcal{P}_{\texttt{C}}$ and $\mathcal{P}_{\texttt{R}}$’s quality. If they are judged as similar-quality, we can be more confident that $\mathcal{P}_{\texttt{R}}$ has needs as valid as $\mathcal{P}_{\texttt{C}}$. To do this, we 0-shot prompt GPT-4o to judge if $\mathcal{P}_{\texttt{C}}$ or $\mathcal{P}_{\texttt{R}}$ is higher-quality, yielding a persona preference $y$. We shuffle the personas and set $y=\texttt{C}\ \textrm{(}\text{or }\texttt{R}$) if $\mathcal{P}_{\texttt{C}}$ (or $\mathcal{P}_{\texttt{R}}$) win in both orders; otherwise, $y=\texttt{Tie}$.
On all datasets except BeaverTails, GPT-4o rates $\mathcal{P}_{\texttt{C}}$ and $\mathcal{P}_{\texttt{R}}$ as similar in quality (Figure 4); the two persona types have very similar win rates (mean difference of 0.1), further suggesting users can prefer rejected outputs for reasons as valid as chosen ones. The large win rate difference of 0.3 on BeaverTails implies $\mathcal{P}_{\texttt{C}}$ conveys user needs that GPT-4o tends to prefer over the needs in $\mathcal{P}_{\texttt{R}}$; we find these preferences may be linked to verbose responses in section 3.4.
3.3 Humans Rate Personas as High-Quality
To confirm LLM personas are high-quality beyond LLM judgments, one Ph.D. student rates 80 total chosen and rejected personas $\mathcal{P}$ from GPT-4o and L-405B on BeaverTails (randomly shuffled) across four axes. First, they rate Applicability from 1-5: how many users are expected to be represented by $\mathcal{P}$—a proxy for the popularity of reasons found in $\mathcal{P}$. Next, as a sanity check, the student gives binary labels for Plausibility: is the user in $\mathcal{P}$ likely to exist; Harmfulness: if $\mathcal{P}$ describes harmful or unethical traits we do not want to train models on; and Overfitting: if $\mathcal{P}$ directly repeats text in prompts or responses, rather than inferring high-level traits.
Chosen and rejected personas are both plausible users, but rejected personas are less applicable (Figure 5), aligning with our intuition that rejected personas are less popular/common but nevertheless valid user needs (section 2.2). Chosen personas are also rarely harmful or overfit, indicating they can form high-quality personalization training signals (section 4).
@@ 3.3 Humans Rate Personas as High-Quality
Table 1: Persona word saliency. Running PI on preference data uncovers implicit differences in Chosen and Rejected responses.

### v2

@@ Abstract
LLMs are aligned to follow input instructions by learning which of two responses users prefer for a prompt. However, such preference data do not convey why users prefer responses that are chosen or rejected, so LLMs trained on these datasets cannot tailor responses to varied user needs. To surface these parameters of personalization, we apply abductive reasoning to preference data, inferring needs and interests of users, i.e., personas, that may prefer either response. We test this idea in two steps: Persona Inference (PI)—abductively inferring personas of users who prefer chosen or rejected outputs—and Persona Tailoring (PT)—training models to tailor outputs to personas from PI. We show: 1) LLMs infer personas accurately explaining why different users may prefer both chosen or rejected outputs; 2) Training on preference data augmented with PI personas via PT boosts personalization and generalizes to supporting user-written personas; and 3) Rejected response personas form harder personalization evaluations, showing PT better aids users with uncommon preferences versus typical alignment methods. We argue for an abductive view of preferences for personalization, asking not only which response is better but when, why, and for whom. [^1] Code and datasets are available at: https://github.com/Pinafore/alignment-personalization
@@ 1 Every Preference Happens for a Reason
Current methods for aligning large language models (LLMs) predominantly use preference data Ji et al. (2023), created by finding: for a given prompt, which of two outputs do users prefer? LLMs are then trained on this data via preference tuning methods like direct preference optimization (Rafailov et al., 2024, DPO), learning to give outputs like the majority chosen response and unlike the rejected one. This improves LLM outputs in dialogue Köpf et al. (2024), question answering (Fan et al., 2019, qa), and summarization Völske et al. (2017).
While preference datasets are valuable, they assume chosen responses are universally better, failing to consider why users prefer responses Joshi et al. (2025). In reality, some users genuinely prefer rejected outputs, even if their reasons are less common. In Figure 2 (left), most users prefer the chosen output for taking brownies to a cake sale. While users valuing simplicity may prefer this direct response, practical users may favor the rejected one, as it also gives packaging logistics (Figure 2, middle). Since users prefer responses for varied reasons Kirk et al. (2024), models trained on preference data should personalize outputs to meet these specific, individual needs Salemi et al. (2024).
A common, interpretable method for personalization involves users specifying their needs and interests as an extra input (e.g., “I like short answers”)—forming a system prompt persona to tailor model responses Zhang et al. (2024c). However, current preference dataset formats lack personas explaining why users prefer responses, overlooking signals for training personalized models Lee et al. (2024).
Figure 1: Training methods with typical preference datasets like DPO cannot fully cater to a user’s specified personas. To overcome this, we train models on preference data augmented with LLM-inferred personas, which we call persona tailoring.
@@
The only way for models trained on existing preference dataset formats to use a persona is by adding it to the inference prompt, hoping the model treats it as an extra instruction to follow Deb et al. (2022). However, standard preference tuning methods like DPO cannot naturally adapt to personas in this way. In Fig 1, DPO tells a sober user to hire a bartender (left), while a user desiring a small list gets 10 ideas (right); such responses are unhelpful, inappropriate, and do not tailor to the users’ specified personas.
As models’ poor personalization stems from a lack of training on personas, we propose abductive reasoning Peirce (1974) to augment preference training data with LLM-inferred personas. Abduction infers hidden contexts that explain a given outcome Zhao et al. (2024). Similarly, we adopt this reasoning to infer (i.e., adduce) hidden needs, interests, and traits of users, i.e., personas Tseng et al. (2024), that explain why chosen and rejected outputs may be preferred. If LLMs surface valid personas of users who may prefer chosen or rejected responses, we can attach them to preference datasets and train models that tailor to user-specified needs, enhancing model personalization (Figure 1, blue).
@@
We first test LLMs in PI on dialogue Bai et al. (2022a), qa Ji et al. (2024), and education Balepur et al. (2024c) preference data. LLM personas accurately convey different users who could prefer chosen or rejected outputs; LLaMA-405B has 91% accuracy, judged by GPT-4o with 90% human agreement (section 3.1). Further, chosen and rejected response personas are often judged as tied in quality (section 3.2), and humans rate rejected ones as plausible but applying to fewer users (section 3.3). Thus, users may prefer rejected outputs for uncommon but valid reasons. Personas are also a useful content analysis tool to find differences in chosen and rejected responses; in BeaverTails Ji et al. (2024), chosen response personas describe users who are “meticulous”, while rejected ones describe “direct” users, showing these labelers may prefer verbosity Zheng et al. (2024a).
As PI yields accurate personas, we augment preference data with PI personas using LLaMA-405B. We then train LLaMA-8B on this new data for the reverse task of PT: using prompts and inferred personas as inputs to give tailored responses (Figure 2, right). We test three strategies: prompting (Brown et al., 2020, $\textsc{PT}_{\textsc{fs}}$), fine-tuning (Chung et al., 2024, $\textsc{PT}_{\textsc{sft}}$), and DPO (Rafailov et al., 2024, $\textsc{PT}_{\textsc{dpo}}$).
Each generation strategy largely boosts personalization when using PT (section 5.1), with $\textsc{PT}_{\textsc{dpo}}$ being the strongest (section 5.2). Further, DPO fine-tuned on preference datasets without personas cannot always tailor to personas during inference; notably, $\textsc{PT}_{\textsc{dpo}}$ is judged as much stronger than DPO on uncommon but still valid needs in the personas from rejected responses (66% average improvement in personalization), showing rejected responses can form valuable, harder evaluations for personalization (section 5.3). Finally, eight users author 144 diverse personas and rate $\textsc{PT}_{\textsc{dpo}}$ and DPO responses for these personas (section 5.4). The same users find our $\textsc{PT}_{\textsc{dpo}}$ method personalizes more effectively to their written needs, showing models trained on realistic, LLM-inferred personas can generalize to real user-specific needs.
@@
2) We release persona-augmented question answering, dialogue, and education preference datasets.1
3) Persona tailoring—prompting and training on persona-augmented data—effectively personalizes to needs inferred by PI and specified by real users.
@@ 2 Persona Inference Setup
The first step in our personalization method is inferring why users may prefer each response in standard preference data (Fig 2, middle). We use abduction Peirce (1974)—which explains outcomes—to infer persona-based explanations for why responses may be preferred via Persona Inference (PI):
@@
Following Chen et al. (2024a), $\mathcal{P}_{1}$ is closest to a demographic persona, but we infer broad traits—like information needs, interests, or personalities—rather than protected attributes (e.g., race) to curb stereotyping Kantharuban et al. (2024). Our personas include no other constraints. All $\mathcal{P}_{1}$ follow the format: “The user is [attribute] and prefers [explanation of preference]” for parsing. This section gives our PI models (section 2.1) and datasets (section 2.2).
@@ 2.1 Models
We test nine LLMs in PI: Claude (Anthropic, 2023, Sonnet, Haiku, Opus), GPT (Achiam et al., 2023, 3.5, 4, 4o), and LLaMA-3.1 Instruct (Dubey et al., 2024, 8B, 70B, 405B). We use 5-shot prompts in the format below, where highlights are generations:
@@ 2.2 Datasets
1) BeaverTails Ji et al. (2024) has advice queries $p$ and candidate answers $\mathcal{R}$ on 14 harm categories including politics, privacy, and unethical behavior.
2) Stanford Human Preferences (Ethayarajh et al., 2022, SHP) has Reddit post questions $p$ on r/ask or advice forums with user-written answers $\mathcal{R}$.
@@
4) Mnemonic Balepur et al. (2024c) has vocabulary terms $p$ and keyword mnemonics $\mathcal{R}$. Mnemonics are study aids that help users learn $p$’s meaning.
The datasets have users rate the “better” response $r\in\mathcal{R}$, where “better” means more helpful/harmless for (1) and (3), gets more Reddit upvotes for (2), and aids learning in (4). The overall better $r$ is chosen ($r_{\texttt{C}}$) and the other is rejected ($r_{\texttt{R}}$). For each entry, we alter the inputs $r_{1}$ and $r_{2}$ in PI to get chosen personas $\mathcal{P}_{\texttt{C}}=\textsc{PI}(p,r_{1}=r_{\texttt{C}},r_{2}=r_{\texttt{R}})$ and rejected personas [^2] We use “rejected” for brevity, not to imply they are worse. $\mathcal{P}_{\texttt{R}}=\textsc{PI}(p,r_{1}=r_{\texttt{R}},r_{2}=r_{\texttt{C}})$ for $r_{\texttt{C}}$ and $r_{\texttt{R}}$. As $r_{\texttt{R}}$ is preferred less often, we assume $\mathcal{P}_{\texttt{R}}$ has less common/popular needs (section 3.3).
In BeaverTails and SHP, some responses $r_{\texttt{R}}$ are deliberately low-quality, with harmful or inaccurate text Liu et al. (2024b). These are out-of-scope, as all $\mathcal{P}_{\texttt{R}}$ are unsafe [^3] Such as “the user is unethical” or “likes misleading users”. and we do not want models to tailor to them (section 9). Thus, we use the data split labeled safe in BeaverTails and outputs with 10+ upvotes in SHP. Post-filtering, we sample 300 entries in each dataset to form 600 PI inputs (Appendix A.1).
3 Evaluating LLM-Inferred Personas
@@
We first verify LLM persona quality before using them to train personalized models (section 4). As many personas can explain the same preference, [^4] For the prompt “What should I eat tonight?” with responses “Pizza!” and “Go eat steak”, a user may prefer the former if they like vegetarian options, enthusiasm, or brevity. we lack ground truth. Instead, we verify that personas accurately explain why users may prefer preference data responses (section 3.1)—the goal of PI (section 2) and a common abduction metric Balepur et al. (2025a)—with GPT-4o (90% human agreement). We then further study personas, showing chosen and rejected personas are similarly valid needs (section 3.2, section 3.3) and reveal preference dataset trends (section 3.4). Thus, PI yields high-quality personas we can use in PT (section 4).
@@ 3.1 LLMs Accurately Infer Personas
We first verify each LLM persona $\mathcal{P}_{1}$ fulfills abduction’s goal: accurately justifying why its response $r_{1}$ may be favored over $r_{2}$ (section 2). GPT-4o—a judge with 90% agreement with three Ph.D. students (Appendix A.4)—evaluates this, judging if a user described by $\mathcal{P}_{1}$ would prefer response $r_{1}$ or $r_{2}$ for the prompt $p$. A chosen persona $\mathcal{P_{\texttt{C}}}$ is accurate if GPT-4o selects $r_{\texttt{C}}$ over $r_{\texttt{R}}$ (and vice versa for $\mathcal{P_{\texttt{R}}}$).
LLMs infer accurate $\mathcal{P_{\texttt{C}}}$ and $\mathcal{P_{\texttt{R}}}$; the judge often picks the intended output (Figure 3). $\mathcal{P_{\texttt{R}}}$ is usually less accurate, consistent with work showing LLMs struggle to justify incorrect answers Balepur et al. (2024a). However, some models show small gaps in accuracy (0.06 for L-405B), so while $\mathcal{P_{\texttt{R}}}$ is harder to infer, it can still reveal plausible needs of users. Finally, accuracy on Mnemonic is lowest, as LLMs must infer why outputs aid learning, which is likely harder than why they are helpful or harmless. Thus, in specific domains Padmakumar et al. (2024), researchers may need to directly elicit why users prefer responses during preference collection for improved accuracy, versus inferring them with LLMs.
@@ 3.2 LLMs Judge Personas as Similar Quality
As LLMs infer accurate chosen/rejected personas $\mathcal{P}_{\texttt{C}}$/$\mathcal{P}_{\texttt{R}}$, we now compare $\mathcal{P}_{\texttt{C}}$ and $\mathcal{P}_{\texttt{R}}$’s quality. If they are judged as similar-quality, we can be more confident that $\mathcal{P}_{\texttt{R}}$ has needs as valid as $\mathcal{P}_{\texttt{C}}$. Thus, we zero-shot prompt GPT-4o to judge if $\mathcal{P}_{\texttt{C}}$ or $\mathcal{P}_{\texttt{R}}$ is higher-quality, yielding a persona preference $y$. We shuffle the personas and set $y=\texttt{C}\ \textrm{(}\text{or }\texttt{R}$) if $\mathcal{P}_{\texttt{C}}$ (or $\mathcal{P}_{\texttt{R}}$) win in both orders; otherwise, $y=\texttt{Tie}$.
On all datasets except BeaverTails, GPT-4o rates $\mathcal{P}_{\texttt{C}}$ and $\mathcal{P}_{\texttt{R}}$ as similar in quality (Figure 4); the two persona types have very similar win rates (mean difference of 0.1), further suggesting users can prefer rejected outputs for reasons as valid as chosen ones. The large win rate difference of 0.3 on BeaverTails implies $\mathcal{P}_{\texttt{C}}$ conveys user needs that GPT-4o tends to prefer over the needs in $\mathcal{P}_{\texttt{R}}$; these preferences may be associated with verbose responses (section 3.4).
@@
Table 1: The top-3 most salient tokens uncovered in personas inferred from chosen and rejected responses. Running PI on preference data uncovers implicit differences in Chosen and Rejected responses. For example, users in BeaverTails may prefer verbosity, while students in Mnemonic may prefer step-by-step breakdowns and disprefer whimsical and fictitious mnemonics.
3.3 Humans Rate Personas as High-Quality
To ensure LLM personas are high-quality beyond LLM judges, three Ph.D. students (Appendix A.12) rate 80 total chosen and rejected personas $\mathcal{P}$ from GPT-4o/L-405B on BeaverTails (randomly shuffled) on four axes. They first rate Applicability on 1–5: how many users are expected to be described by $\mathcal{P}$—a proxy for the popularity of reasons in $\mathcal{P}$. Next, as a sanity check, the students provide binary labels for Plausibility: is the user in $\mathcal{P}$ likely to exist; Harmfulness: if $\mathcal{P}$ describes harmful or unethical traits we do not want to train models on; and Overfitting: if $\mathcal{P}$ directly repeats text in prompts or responses, rather than inferring high-level traits.
Chosen and rejected personas are both plausible users, further confirming their accuracy (section 3.1), but rejected ones are slightly less applicable (Figure 5), supporting the intuition that rejected personas are less common but valid needs (section 2.2). Chosen personas are rarely harmful or overfit, so they can form high-quality training data for personalization (section 4).